OpenStudy (anonymous):

if f(x) = x − 4 and g(x) = 1/2 x − 2, find f[g(−6)].

OpenStudy (stamp):

f(g(-6)) So find g(-6), and plug it in f( ).

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (stamp):

g(-6) = 1/2 (6) - 2

OpenStudy (stamp):

See how I plugged in 6? So now take whatever you find g(-6) is equal to and plug it into f(x). f(g(-6))= g(-6) - 4

OpenStudy (anonymous):

i got 1/3 thats not one of the options

OpenStudy (anonymous):

9 3 −9 −3

OpenStudy (anonymous):

those are my options so i did it wrong

OpenStudy (anonymous):

please i need the answer

OpenStudy (stamp):

if f(x) = x − 4 and g(x) = 1/2 x − 2, find f[g(−6)]. g(-6) = 1/2 (-6) - 2 = (-3) - 2 = -5 g(-6) = -5 f(g(-6)) = (g(-6)) - 4 f(g(-6)) = (-5) - 4 f(g(-6)) = -9